Exam By MH Botros 1.A car with a mass of 1200 kg is moving around acu with a radius of m constant speed of 10 m/s Calculate the centripetal acceleration the the two decimal the final answer Determine the magnitude of the force required to produce this centri acceleration. Round to nearest whole number 2. A 0.23 kg ball moving in a circle at the end of a string has a centripe acceleration of 10 m/s. Determine the magnitude of the centripetal force exerted by the string on the ball to produce this acceleration. Round the final answer to one decimal place. 3. Two masses are attracted by a gravitational force of0.12 N. What will the force of attraction be if the distance between the two masses is halved? Round the final answer to two decimal places. 4- Two 600-kg masses (1332 are separated by a distance of64 m. Using Newtons law oferavitation,find Round the final answer to two decimal places. s. At a given instant in time, a 10-kg rock is observed to be falling with an acceleration of 8.5 m/s. What is the magnitude of the force of air resistance acting upon the rock at this instant? Round the final answer to the nearest whole number. 6. A single force of60 N acts upon a 14-kg block. What is the magnitude of the acceleration of the block? Round the final answer to the nearest whole number. 7. A 8-kg block being pushed across a table by a force Pacting in the positive direction has an acceleration of 2 m/s also in the positive direction. If the magnitude of P is 30 N, what is the magnitude of the frictional force acting upon the block? Round the final answer to the nearest whole number 8.A heavy ball with a mass of 12.5 kg is observed to at a rate of28 m/s. What is the size of the net force acting on this ball? Round the final answer to the nearest whole number 0.8 kg book rests on a table. A downward force of 11 Nis exerted on the top of the book by a hand pushing down the book. force acting upon the book? Round the final answer to two What is the magnitude of the gravitational decimal places. upward (normal) force exerted by the table on the book? Round the final b. What is the magnitude of the answer to two decimal places c. Is the book accelerated? At a given instant in time, a 5-kg rock that has been dropped from a high cliffexperiences a force of air resistance of 15 N. What are the magnitude and direction of the acceleration of the rock? (Do not forget the gravitational force!) Round the final answer to two decimal places. Extracedit 11. You stand on a scale that rests on the floor of an elevator that is accelerating upward. What is the relationship between the force due to gravity and the normal force exerted by the scale A. N mg D. Na 0 C. NKmg
Answer #1

1) a) Centripetal acceleration ac = v2 / r = 102 / 35

= 2.86 m/s2

b) from Newton’s 2nd Law of motion, F = m a

= 1200 * 2.86

F = 3432 kg.m/s2